# Spatiotemporal-Augmented Graph Neural Networks for Human Mobility Simulation

### v1

@@ 5.7 RQ4: Visualization
The geographical visualization results of our proposed STAR method and two baseline methods on the TKY dataset are presented in Figure 5. The side colorbar illustrates the frequency of location visits, with brighter colors indicating higher visit frequencies. The STAR method yields more similar results in the diagonal direction’s highlighted areas to the real visit frequencies than the other two methods, namely CGE and IO-HMM, which are one of state-of-the-art deep learning methods and statistical modeling methods for trajectory simulation, respectively. Additionally, STAR also provides attention to lower-visited areas in the upper-left corner, while the other two methods lack precision in modeling the long-tailed location visit patterns.
@@ 6 Conclusion
In this paper, we propose a novel framework for human mobility simulation, namely SpatioTemporal-Augmented gRaph neural networks (STAR), to model the spatiotemporal effects among locations. On the one hand, STAR designs various kinds of spatiotemporal graphs to incorporate the spatiotemporal correspondences of locations, revealing the spatial proximity and functionality similarity with the temporal visit distribution. On the other hand, STAR builds the dual-branch decision generator module to balance the diverse transitions generated by the exploration branch and the repetitive patterns generated by the dwell branch. The STAR framework is optimized based on the classification rewards of the policy discriminator module after iteratively generating a complete trajectory sequence. We have conducted comprehensive experiments on the human mobility simulation task, verifying the superiority of the STAR framework to the state-of-the-art methods. Ablation studies on the multi-channel embedding module reveal that different datasets prefer different spatiotemporal graphs. Elaborate experiments further verify the effectiveness of the edges of the Vanilla version and the dwell branch.
However, deep learning-based models always suffer from data scarcity and sparsity, easily leading to overfitting in the small training data. In the future, we will explore the commonalities of different human mobility data and attempt to transfer the human mobility model from high-resource locations to low-resource locations. Another interesting direction is introducing massive external resources like daily social tweets to build a huge location graph, which can largely alleviate the data sparsity problem.

### v3

@@ 5.7 RQ4: Visualization
We visualize the location visit frequency of all methods on the four datasets to distinguish their simulated trajectories from the preference perspective (i.e. G-rank metric). Due to space limitations, we present the results of the real data and three representative methods on the first dataset (NYC) in Figure 5. It can be observed that the location visit frequency scale of STAR is closer to the real data compared with the other two baseline methods. However, the visited locations in trajectories simulated by CGE are more concentrated, where some highly visited locations are even not visited in the real data. Conversely, IO-HMM exhibits more attentions on the lower-visited areas, resulting in a relatively scattered distribution of visited locations.
@@ 6 Conclusion
In this paper, we propose a novel framework for human mobility simulation, namely SpatioTemporal-Augmented gRaph neural networks (STAR), to model the spatiotemporal effects among locations. On the one hand, STAR designs various kinds of spatiotemporal graphs to incorporate the spatiotemporal correspondences of locations, revealing the spatial proximity and functional similarity with the temporal visit distribution. On the other hand, STAR builds the dual-branch decision generator module to balance the diverse transitions generated by the exploration branch and the repetitive patterns generated by the dwell branch. The STAR framework is optimized based on the classification rewards of the policy discriminator module after iteratively generating a complete trajectory sequence. We have conducted comprehensive experiments on the human mobility simulation task, verifying the superiority of the STAR framework to the state-of-the-art methods. Ablation studies on the multi-channel embedding module reveal that different datasets prefer different spatiotemporal graphs. Elaborate experiments further verify the effectiveness of the edges of the Vanilla version and the dwell branch.
Limitations and future works. The limitations of our work stem from the lack of dynamic scenarios and data scarcity. Firstly, our proposed STAR lacks consideration for dynamic scenarios, such as massive commuting patterns and traffic accidents, which may result in an inaccurate capture of real-world human mobility behaviors and limit its applicability. Secondly, similar to other deep learning methods, our proposed STAR may encounter performance degradation in scenarios with limited data availability for human mobility simulation. In future work, we plan to consider more external factors as well as mobility commonalities of different human mobility data to comprehensively enhance the mobility simulation task.
7 Ackowledgement
This research is supported by the Joint Funds of the Zhejiang Provincial Natural Science Foundation of China under (Grant No. LHZSD24F020001), Zhejiang Province ”LingYan” Research and Development Plan Project (No. 2024C01114).